Project

General

Profile

Actions

Ops/Comm Scheduling meeting : 14/11/2017

Location : CR21 at 14:00 pm

Phone access : phonebridge number
UK: 08082380274
US: 631.621.5253
Access code: 1779339517

  • Actions:
    • Need to give Tony Millington heads up on survey requirements in Nov/Dec once the run plan has settled (3 Absorber changes)
  • DC Update
  • Report on pulser run
  • Report on SSU cryo 5
  • Shutdown activities - the list so far
Name Task Time required Done?
Mark Partial warming of DS to clean vacuum space and recool 10 days Cooling
Ajit/MOM Review current alarms 1 day Went through alarms with CM, all understood, some state changes req'd
Joe/Henry Change target patch fibres 0.5 days In progress, Joe down Tues/Weds. Inspect in synch
John/Colin Hall AC work? ? Proceeding.
Turn on tracker helium Melissa and Durga
Finalise tracker calibration Melissa
  • Next user cycle : 2017/03 : Starts today / shifts start tomorrow
    • Shifts are now open.
Shift Block # Open shifts / 6 # data-taking hours / day
1 0.8 16 No data-taking during weekday morning shifts; non-critical gaps on weekend.
2 2 16 No data-taking during weekday morning shift / Possible absorber change
3 1 16 No data-taking during weekday morning shift / Possible absorber change
4 0 24 Magnet change during day shift
5 0 24 Magnet change during day shift; Possible absorber change
6 2 16 No data-taking during weekday; Possible absorber change
7 0 24 Magnet change during day shift
Shifts for the weekend:
Night Morning Evening
Saturday Milorad/Paul K Kevin/? Ray / Chris
Sunday Milorad/Paul K Kevin/Ray Chris / ?
  • Need one person on Saturday morning and one person on Sunday evening shift to permit full data-taking over the weekend.
  • Can support some data-taking during shift block 3, but not every day.
  • Plan to start data-taking.
    • Run plan : attachment:2017-03-RunPlan-v5.xlsx
      • Fallback run plan : attachment:2017-03-RunPlan-v6.xlsx
  • AOB

Updated by Whyte, Colin almost 6 years ago · 10 revisions